# Semantically Diverse Language Generation for Uncertainty Estimation in Language Models

### v1

@@ Abstract
Large language models (LLMs) can suffer from hallucinations when generating text. These hallucinations impede various applications in society and industry by making LLMs untrustworthy. Current LLMs generate text in an autoregressive fashion by predicting and appending text tokens. When an LLM is uncertain about the semantic meaning of the next tokens to generate, it is likely to start hallucinating. Thus, it has been suggested that hallucinations stem from predictive uncertainty. We introduce Semantically Diverse Language Generation (SDLG) to quantify predictive uncertainty in LLMs. SDLG steers the LLM to generate semantically diverse yet likely alternatives for an initially generated text. This approach provides a precise measure of aleatoric semantic uncertainty, detecting whether the initial text is likely to be hallucinated. Experiments on question-answering tasks demonstrate that SDLG consistently outperforms existing methods while being the most computationally efficient, setting a new standard for uncertainty estimation in LLMs.
@@ 1 Introduction
Hallucinations hinder a broad use of LLMs in practical applications and critical decision-making processes as they make them untrustworthy (Manakul et al., 2023). Hallucinations are fragments of generated text that, despite appearing cohesive, are not factual. We consider generated text to be hallucinated if it stems from contradictory or non-existent facts in the training data or input prompt. Thus, hallucinations are conjectured to be mainly caused by the predictive uncertainty inherent to probabilistic models (Xiao and Wang, 2021). While uncertainty estimation for classification tasks has been developed extensively (Hüllermeier and Waegeman, 2021; Gawlikowski et al., 2023), it remains a challenging problem for autoregressive tasks, particularly in natural language generation (NLG).
@@
Figure 1: Standard multinomial sampling relies on chance to obtain semantically different output sequences, thus is prone to miss them. SDLG specifically searches for likely, but semantically different output sequences, thereby improving the estimation of semantic uncertainty in language models.
To improve upon the limitations of the current state-of-the-art approach, we introduce Semantically Diverse Language Generation (SDLG), a method that efficiently estimates semantic uncertainty by utilizing importance sampling to generate output sequences. We introduce a proposal distribution that samples semantically diverse output sequences (see Fig. 1). SDLG utilizes the NLI model not only for transforming the space of generated output sequences to semantic clusters but also for computing the contribution of every token to the final semantics. Subsequently, the semantically most relevant tokens are substituted to explicitly steer the generation toward semantically diverse yet likely alternative output sequences, while correcting for the increased sampling probability using importance sampling. This can be viewed as stress testing the language model, unveiling output sequences that are valuable summands for estimators of semantic uncertainty. They are generated by the current sampling methods only by chance, making SDLG a more systematic and reliable method for capturing semantic uncertainty.
Our main contributions are:
  • We propose a novel method for generating semantically diverse yet likely output sequences. Empirical results demonstrate that our method outperforms existing methods for uncertainty estimation in NLG, specifically across a variety of free-form question-answering tasks.
  • We establish a theoretical foundation for uncertainty measures in NLG and introduce theoretically grounded estimators for aleatoric semantic uncertainty, also known as semantic entropy. Applying these estimators enhances the empirical performance of uncertainty estimation in language models.
@@ 2 Measuring Predictive Uncertainty in NLG
Uncertainty estimation in classification tasks has already been well-established (Gal and Ghahramani, 2016). However, these measures cannot be directly applied to uncertainty estimation in NLG. Two key aspects, which differ from estimating uncertainty in classification tasks, have to be considered. First, a sequence of autoregressive predictions collectively forms the final output of a model. Second, unlike classification tasks where classes usually are mutually exclusive, different output sequences may be equivalent in their semantic meaning. To account for this aspects, Kuhn et al. (2023) introduce semantic entropy, yet without proper theoretical grounding. In the following, we derive semantic entropy starting from the well-studied uncertainty estimation in classification tasks.
@@
We briefly revisit uncertainty estimation for classification tasks. In this work, we quantify the predictive uncertainty of a single given “off-the-shelf” model. Given are a classification model parametrized by $\bm{w}$ and an input vector $\bm{x}$. The predictive distribution under the given model is denoted as $p(y\mid\bm{x},\bm{w})$. We assume a fixed dataset $\mathcal{D}$ that was sampled according to the predictive distribution $p(y\mid\bm{x},\bm{w}^{*})$ under true model parameters $\bm{w}^{*}$. Thus, we assume that the model class can approximate the true distribution sufficiently well, a common and usually necessary assumption (Hüllermeier and Waegeman, 2021). The posterior distribution $p(\bm{w}\mid\mathcal{D})$ denotes how likely $\bm{w}$ matches $\bm{w}^{*}$. Following Schweighofer et al. (2023a, b), the predictive uncertainty of a single given model parametrized by $\bm{w}$ is given by
@@
is used to derive the predictive uncertainty in the NLG setting. Here, $p(\bm{y}\mid\bm{x},\bm{w})$ expresses the probability of generating an output sequence $\bm{y}$ and $p(c\mid\bm{y},\bm{x},\bm{w})$ expresses the probability of $\bm{y}$ belonging to a certain semantic cluster $c\in\mathcal{C}$. Although a specific output sequence might potentially be attributed to more than one semantic cluster, we follow Kuhn et al. (2023) in assuming that each output sequence is attributed to a single semantic cluster. They demonstrate that assigning semantic clusters by predicting the semantic equivalence between generated output sequences with an NLI model empirically performs well. The NLI model takes two sequences as input and predicts whether they entail or contradict each other. Two output sequences are semantic equivalent if they entail each other in both orders and thus belong to the same semantic cluster by definition. Consequently, $p(c\mid\bm{y},\bm{x},\bm{w})$ is equivalent to $\mathbbm{1}\!\{\bm{y}\in c\mid\bm{x},\bm{w}\}$, where $\mathbbm{1}\!\{\bm{y}\in c\mid\bm{x},\bm{w}\}=1$ iff $\bm{y}$ belongs to semantic cluster $c$. For the set of all possible output sequences $\mathcal{Y}$, $p(c\mid\bm{x},\bm{w})$ expresses the probability of the language model generating an output sequence belonging to a specific semantic cluster for a given input sequence. Adopting the definition of predictive uncertainty in Eq. (1), the total predictive semantic uncertainty
@@
can again be additively decomposed into aleatoric and epistemic semantic uncertainty. The epistemic semantic uncertainty is again a posterior expectation, which is particularly challenging to estimate for current language models with billions of parameters (Zhang et al., 2022; Touvron et al., 2023). The aleatoric semantic uncertainty turns out to be precisely the semantic entropy proposed by Kuhn et al. (2023). Semantic entropy is the entropy of the semantic cluster probability distribution as of Eq. (2),
@@
under a single given language model. To determine whether a language model is uncertain about its output, it is essential to accurately estimate its semantic entropy, as addressed in the following section.
@@ 3 Estimating Aleatoric Semantic Uncertainty in NLG
$p(c\mid\bm{x},\bm{w})$ $\approx\ \frac{1}{N}\ \sum_{n=1}^{N}\mathbbm{1}\!\{\bm{y}\in c\mid\bm{x},\bm{w}\}\ ,\quad\bm{y}^{n}\sim p(\bm{y}\mid\bm{x},\bm{w})\ .$ (6)
Output sequences $\bm{y}^{n}$ are simply generated via multinomial sampling (Kuhn et al., 2023). This estimate of the semantic cluster probability distribution can directly be used to approximate the semantic entropy. Since usually not all clusters $c\in\mathcal{C}$ are found through sampling, the sum is taken over all observed clusters $c_{1},...,c_{M}$ to which $\{\bm{y}^{n}\}_{n=1}^{N}$ were assigned by the NLI model, resulting in
@@
In Sec. 6 we empirically show that using this proper estimator for the semantic entropy outperforms the improper estimator implemented by Kuhn et al. (2023). For further details see Sec. C in the appendix.
@@
$p(c\mid\bm{x},\bm{w})$ $\approx\ \frac{1}{N}\ \sum_{n=1}^{N}\mathbbm{1}\!\{\bm{y}\in c\mid\bm{x},\bm{w}\}\ \frac{p(\bm{y}^{n}\mid\bm{x},\bm{w})}{q(\bm{y}^{n}\mid\bm{x},\bm{w})}\ ,\quad\bm{y}^{n}\sim q(\bm{y}\mid\bm{x},\bm{w})\ .$ (8)
@@ 4 Semantically Diverse Language Generation
Estimating the semantic entropy according to Eq. (7) requires approximating the semantic cluster probability distribution. This probability distribution can either be approximated according to Eq. (6), or via importance sampling according to Eq. (8). MC estimation using output sequences from multinomial sampling is straightforward, as one can directly sample from the original distribution $p(\bm{y}\mid\bm{x},\bm{w})$ without the need for additional weighting or adjustment. However, the resulting estimator has high variance and sampling only considers the likelihood and not the semantic diversity. Furthermore, multinomial sampling may miss likely output sequences that capture important information about semantic cluster probabilities, especially if those sequences require choosing a lower-likelihood token. This limits the accurate estimation of the semantic entropy, as sampling semantically diverse output sequences essentially occurs by chance. Importance sampling, on the other hand, can overcome these limitations by incorporating semantic diversity into the sampling procedure. However, this requires a beneficial proposal distribution $q(\bm{y}\mid\bm{x},\bm{w})$. We propose Semantically Diverse Language Generation (SDLG) to sample according to such an empirical proposal distribution. It seeks to efficiently explore semantic clusters, capturing important modes of $p(c\mid\bm{x},\bm{w})$ that might be missed by multinomial sampling.
@@
Given an input sequence $\bm{x}$ and an initial output sequence $\bm{y}^{\prime}$ generated by a language model, how can we generate another output sequence that has a different semantics from $\bm{y}^{\prime}$? In natural language, individual words contribute to the semantics of the sentence to varying extents. For instance, consider the sentence “Albert Einstein proposed the theory of relativity”. The word “proposed” can be substituted with “introduced” or “postulated” without altering the semantics. However, replacing “Einstein” with “Schweitzer” would change the semantics. Therefore, our method focuses on identifying and substituting the tokens that are most critical to the semantics of $\bm{y}^{\prime}$ (see Fig. 1). This is achieved by introducing three scores at the token level that quantify each token’s relevance in altering the semantics:
@@
The computation of the attribution and substitution scores requires a loss $\mathrm{L}$, which expresses to what degree $\bm{y}^{\prime}$ is semantically different from itself. It is computed utilizing an NLI model that predicts whether two sequences entail or contradict each other. First, the initial output sequence is input into the NLI model twice, resulting in a high probability of entailment. Second, the loss $\mathrm{L}$ is computed to the target prediction contradiction. This loss is used to compute the gradient $\nabla_{\bm{z}_{i}}\mathrm{L}$ with respect to the token embedding $\bm{z}_{i}$ that represents the initial token $y_{i}$. This gradient vector quantifies the required change in $\bm{z}_{i}$ to achieve a high probability of contradiction, thus altering the semantic meaning of $\bm{y}^{\prime}$.
@@
$A_{i}$ $=\ \|\bm{z}_{i}\odot\nabla_{\bm{z}_{i}}\mathrm{L}\|_{2}$ (9)
is defined as the Euclidean distance $\|\cdot\|_{2}$ of the gradient vector multiplied elementwise with the embedding vector $\bm{z}_{i}$ that represents the initial token $y_{i}$. The higher the attribution score $A_{i}$, the higher the impact of the token $y_{i}$ on altering the semantics when being changed (Adebayo et al., 2018). We note that different attribution methods could be utilized and future work may benefit from exploring these methods to compute $A_{i}$ (Madsen et al., 2022).
@@
$S_{ij}$ $=\ \frac{(\bm{z}_{i}-\bm{z}_{j})\cdot\nabla_{\bm{z}_{i}}\mathrm{L}}{\|\bm{z}_{i}-\bm{z}_{j}\|_{2}\ \ \|\nabla_{\bm{z}_{i}}\mathrm{L}\|_{2}}$ (10)
@@
Identifying which alternative token should substitute which initial token is important but not sufficient. Not every alternative token $v_{j}$ is an appropriate substitution for the initial token $y_{i}$ given the context. The context is the input sequence $\bm{x}$ and the output sequence up to the token that is to be substituted $\bm{y}_{<i}$. Substituting with a very unlikely token might undermine the coherence of the overall sequence. Thus, our third objective is to favor alternative tokens that exhibit a high likelihood. The alternative token’s importance score
$I_{ij}$ $=\ p(v_{j}\mid\bm{y}_{<i},\bm{x},\bm{w})$ (11)
@@
0:  Language model $g(\cdot)$, input sequence $\bm{x}$, vocabulary $\mathcal{V}$, number output sequences $N$
1:  $\mathcal{S}\leftarrow\emptyset$
2:  $\bm{y}^{1}\leftarrow g(\bm{x})$
3:  $\mathcal{S}\leftarrow\mathcal{S}\cup\{\bm{y}^{1}\}$
@@
1:  $\mathcal{R}\leftarrow\emptyset$
2:  $\mathrm{L}\leftarrow l(e(\bm{y},\bm{y}),c_{\text{contradiction}})$
@@
All token pairs ($y_{i}$, $v_{j}$) are ranked according to the three scores ($A_{i}$, $S_{ij}$, $I_{ij}$). In its simplest form, the ranking is based on equally weighting the three scores, which we found to work well empirically. Based on this token score ranking, a new output sequence is generated by deliberately substituting the highest-ranked token pair. Subsequent tokens are discarded as they are conditioned on the substituted token, which affects their likelihood. The remainder of the new output sequence is then generated by the language model using the usual sampling strategy (see Alg. 1 and 2). SDLG preserves the semantically less relevant part of the output sequence, eliminating the need to regenerate it and focusing on the part with a high chance of altering the semantic meaning. As a result, the generation process becomes computationally more efficient than sampling from scratch each time and potentially sampling multiple duplicate output sequences.
@@
Finally, we can discuss the exact form of the proposal distribution $q(\bm{y}\mid\bm{x},\bm{w})$ in Eq. (8), that is induced by SDLG. We define the proposal distribution as
@@
where $p(y_{t}\mid\bm{y}_{<t},\bm{x},\bm{w})$ is the likelihood of the token that is exchanged by SDLG. Intuitively, it means that we have to adjust the MC estimate in Eq. (6), because SDLG interferes in sampling and changes the token $y_{t}$ deterministically. For more details on the assumptions and a step-by-step derivation see Sec. B in the appendix. Our experiments show, that sampling according to this proposal distribution improves uncertainty estimation in NLG.

### v2

# Improving Uncertainty Estimation through Semantically Diverse Language Generation

@@ Abstract
Large language models (LLMs) can suffer from hallucinations when generating text. These hallucinations impede various applications in society and industry by making LLMs untrustworthy. Current LLMs generate text in an autoregressive fashion by predicting and appending text tokens. When an LLM is uncertain about the semantic meaning of the next tokens to generate, it is likely to start hallucinating. Thus, it has been suggested that predictive uncertainty is one of the main causes of hallucinations. We introduce Semantically Diverse Language Generation (SDLG) to quantify predictive uncertainty in LLMs. SDLG steers the LLM to generate semantically diverse yet likely alternatives for an initially generated text. This approach provides a precise measure of aleatoric semantic uncertainty, detecting whether the initial text is likely to be hallucinated. Experiments on question-answering tasks demonstrate that SDLG consistently outperforms existing methods while being the most computationally efficient, setting a new standard for uncertainty estimation in LLMs.
@@ 1 Introduction
Hallucinations hinder a broad use of LLMs in practical applications and critical decision-making processes as they make them untrustworthy (Manakul et al., 2023). Hallucinations are fragments of generated text that, despite appearing cohesive, are not factual. At the time of writing, there is no consensus on the exact nature of all causes of hallucination. We consider generated text to be hallucinated if it stems from contradictory or non-existent facts in the training data or input prompt. Such hallucinations are conjectured to be mainly caused by the predictive uncertainty inherent to probabilistic models (Xiao and Wang, 2021). This type of hallucination is also referred to as confabulation (Farquhar et al., 2024), and we shall use the two terms interchangeably. While uncertainty estimation for classification tasks has been developed extensively (Hüllermeier and Waegeman, 2021; Gawlikowski et al., 2023), it remains a challenging problem for autoregressive tasks, particularly in natural language generation (NLG).
@@
Figure 1: Using standard multinomial sampling to generate text does not account for its semantics. Thus, it relies on chance to obtain semantically diverse output sequences and is prone to miss them. SDLG addresses this by specifically searching for likely, but semantically different output sequences. Thereby, the estimation of semantic uncertainty in language models is improved.
To improve upon the limitations of the current state-of-the-art approach, we introduce Semantically Diverse Language Generation (SDLG), a method that efficiently estimates semantic uncertainty by utilizing importance sampling to generate output sequences. We introduce a proposal distribution that samples semantically diverse output sequences (see Fig. 1). SDLG utilizes the NLI model not only for transforming the space of generated output sequences to semantic clusters but also for computing the contribution of every token to the final semantics. Subsequently, the semantically most relevant tokens are substituted to explicitly steer the generation toward semantically diverse yet likely alternative output sequences, while correcting for the increased sampling probability using importance sampling. This can be viewed as stress-testing the language model, unveiling output sequences that are valuable summands for estimators of semantic uncertainty. They are generated by the current sampling methods only by chance, making SDLG a more systematic and reliable method for capturing semantic uncertainty. Our main contributions are:
  • We propose a novel method for generating semantically diverse yet likely output sequences. Empirical results demonstrate that our method outperforms existing methods for uncertainty estimation in NLG, specifically across a variety of free-form question answering tasks.
  • We establish a theoretical foundation for uncertainty measures in NLG and introduce theoretically grounded estimators for aleatoric semantic uncertainty, also known as semantic entropy. Applying these estimators enhances empirical performance of uncertainty estimation in language models.
@@ 2 Measuring Predictive Uncertainty in NLG
Uncertainty estimation in classification tasks has already been well-established (Gal and Ghahramani, 2016). However, these measures cannot be directly applied to uncertainty estimation in NLG. Two key aspects, which differ from estimating uncertainty in classification tasks, have to be considered. First, a sequence of autoregressive predictions collectively forms the final output of a model. Second, unlike classification tasks where classes usually are mutually exclusive, different output sequences may be equivalent in their semantic meaning. To account for these aspects, Kuhn et al. (2023) introduce semantic entropy, yet without proper theoretical grounding. In the following, we derive semantic entropy starting from the well-studied uncertainty estimation in classification tasks.
@@
We briefly revisit uncertainty estimation for classification tasks. In this work, we quantify the predictive uncertainty of a single given “off-the-shelf” model. Given are a classification model parametrized by $\bm{w}$ and an input vector $\bm{x}$. The predictive distribution under the given model is denoted as $p(y\mid\bm{x},\bm{w})$. We assume a fixed dataset $\mathcal{D}$ that was sampled according to the predictive distribution $p(y\mid\bm{x},\bm{w}^{*})$ under true model parameters $\bm{w}^{*}$. Thus, we assume that the model class can approximate the true distribution sufficiently well, a common and usually necessary assumption (Hüllermeier and Waegeman, 2021). The posterior distribution $p(\bm{w}\mid\mathcal{D})$ denotes how likely $\bm{w}$ matches $\bm{w}^{*}$. Following Schweighofer et al. (2023a; b), the predictive uncertainty of a single given model parametrized by $\bm{w}$ is given by
@@
is used to derive the predictive uncertainty in the NLG setting. Here, $p(\bm{y}\mid\bm{x},\bm{w})$ expresses the probability of generating an output sequence $\bm{y}$ and $p(c\mid\bm{y},\bm{x},\bm{w})$ expresses the probability of $\bm{y}$ belonging to a certain semantic cluster $c\in\mathcal{C}$. Although a specific output sequence might potentially be attributed to more than one semantic cluster, we follow Kuhn et al. (2023); Farquhar et al. (2024) in assuming that each output sequence is attributed to a single semantic cluster. They demonstrate that assigning semantic clusters by predicting the semantic equivalence between generated output sequences with an NLI model empirically performs well. The NLI model takes two sequences as input and predicts whether they entail or contradict each other. Two output sequences are semantic equivalent if they entail each other in both orders and thus belong to the same semantic cluster by definition. Consequently, $p(c\mid\bm{y},\bm{x},\bm{w})$ is equivalent to $\mathbbm{1}\!\{\bm{y}\in c\mid\bm{x},\bm{w}\}$, where $\mathbbm{1}\!\{\bm{y}\in c\mid\bm{x},\bm{w}\}=1$ iff $\bm{y}$ belongs to semantic cluster $c$. For the set of all possible output sequences $\mathcal{Y}$, $p(c\mid\bm{x},\bm{w})$ expresses the probability of the language model generating an output sequence belonging to a specific semantic cluster for a given input sequence. Adopting the definition of predictive uncertainty in Eq. (1), the total predictive semantic uncertainty
@@
can again be additively decomposed into aleatoric and epistemic semantic uncertainty. The epistemic semantic uncertainty is again a posterior expectation, which is particularly challenging to estimate for current language models with billions of parameters (Zhang et al., 2022; Touvron et al., 2023). The aleatoric semantic uncertainty turns out to be precisely the semantic entropy proposed by Kuhn et al. (2023). Semantic entropy is the entropy of the semantic cluster probability distribution,
@@
under a single given language model. To determine if a language model is uncertain about its output, it is essential to accurately estimate its semantic entropy, as addressed in the following section.
@@ 3 Estimating Aleatoric Semantic Uncertainty in NLG
$p(c\mid\bm{x},\bm{w})\$$\approx\ \frac{1}{N}\ \sum_{n=1}^{N}\mathbbm{1}\!\{\bm{y}\in c\mid\bm{x},\bm{w}\}\ ,\quad\bm{y}^{n}\sim p(\bm{y}\mid\bm{x},\bm{w})\ .$ (6)
Output sequences $\bm{y}^{n}$ are simply generated via multinomial sampling (Kuhn et al., 2023). This estimate of the semantic cluster probability distribution can directly be used to approximate the semantic entropy. Since usually not all clusters $c\in\mathcal{C}$ are found through sampling, the sum is taken over observed clusters $c_{1},...,c_{M}$ to which $\{\bm{y}^{n}\}_{n=1}^{N}$ were assigned by the NLI model, resulting in
@@
In Sec. 6, we empirically show that this proper estimator for semantic entropy outperforms the improper estimator implemented by Kuhn et al. (2023). For further details see Sec. C in the appendix.
@@
$p(c\mid\bm{x},\bm{w})\$$\approx\ \frac{1}{N}\ \sum_{n=1}^{N}\mathbbm{1}\!\{\bm{y}\in c\mid\bm{x},\bm{w}\}\ \frac{p(\bm{y}^{n}\mid\bm{x},\bm{w})}{q(\bm{y}^{n}\mid\bm{x},\bm{w})}\ ,\quad\bm{y}^{n}\sim q(\bm{y}\mid\bm{x},\bm{w})\ .$ (8)
@@ 4 Semantically Diverse Language Generation
Estimating the semantic entropy according to Eq. (7) requires approximating the semantic cluster probability distribution. This probability distribution can either be approximated according to Eq. (6), or via importance sampling according to Eq. (8). MC estimation using output sequences from multinomial sampling is straightforward, as one can directly sample from the original distribution $p(\bm{y}\mid\bm{x},\bm{w})$ without the need for additional weighting or adjustment. However, the resulting estimator has high variance, and sampling only considers the likelihood and not the semantic diversity. Furthermore, multinomial sampling may miss likely output sequences that capture important information about semantic cluster probabilities, especially if those sequences require choosing a lower-likelihood token. This limits the accurate estimation of the semantic entropy, as sampling semantically diverse output sequences essentially occurs by chance. Importance sampling, on the other hand, can overcome these limitations by incorporating semantic diversity into the sampling procedure. However, this requires a beneficial proposal distribution $q(\bm{y}\mid\bm{x},\bm{w})$. We propose Semantically Diverse Language Generation (SDLG) to sample according to such an empirical proposal distribution. It seeks to efficiently explore semantic clusters, capturing important modes of $p(c\mid\bm{x},\bm{w})$ that might be missed by multinomial sampling.
@@
Given an input sequence $\bm{x}$ and an initial output sequence $\bm{y}^{\prime}$ generated by a language model, how can we generate another output sequence that has a different semantics from $\bm{y}^{\prime}$? In natural language, individual words contribute to the semantics of the sentence to varying extents. For instance, consider the sentence “Einstein proposed the theory of relativity”. The word “proposed” can be substituted with “introduced” without altering the semantics. However, replacing “Einstein” with “Newton” would change the semantics. Therefore, our method focuses on identifying and substituting the tokens that are most critical to the semantics of $\bm{y}^{\prime}$ (see Fig. 1). This is achieved by introducing three scores at the token level that quantify each token’s relevance in altering the semantics:
@@
Computing the attribution and substitution scores requires a loss $\mathrm{L}$, which expresses to what degree $\bm{y}^{\prime}$ is semantically different from itself. It is computed utilizing an NLI model that predicts whether two sequences entail or contradict each other. First, the initial output sequence is fed into the NLI model twice, resulting in a high probability of entailment. Second, the loss $\mathrm{L}$ is computed for the target prediction contradiction. This loss is used to compute the gradient $\nabla_{\bm{z}_{i}}\mathrm{L}$ w.r.t. the token embedding $\bm{z}_{i}$ that represents the initial token $y_{i}$. This gradient vector quantifies the required change in $\bm{z}_{i}$ to achieve a high probability of contradiction, thus altering the semantic meaning of $\bm{y}^{\prime}$.
@@
$A_{i}\$$=\ \|\bm{z}_{i}\odot\nabla_{\bm{z}_{i}}\mathrm{L}\|_{2}$ (9)
is defined as the Euclidean distance $\|\cdot\|_{2}$ of the gradient vector multiplied elementwise with the embedding vector $\bm{z}_{i}$ that represents the initial token $y_{i}$. The higher the attribution score $A_{i}$, the higher the impact of the token $y_{i}$ on altering the semantics when being changed (Adebayo et al., 2018). We note that different attribution methods could be utilized, and future work may benefit from exploring these methods to compute $A_{i}$ (Madsen et al., 2022).
@@
$S_{ij}\$$=\ \frac{(\bm{z}_{i}-\bm{z}_{j})\cdot\nabla_{\bm{z}_{i}}\mathrm{L}}{\|\bm{z}_{i}-\bm{z}_{j}\|_{2}\ \ \|\nabla_{\bm{z}_{i}}\mathrm{L}\|_{2}}$ (10)
@@
Identifying which alternative token should substitute which initial token is important but not sufficient. Not every alternative token $v_{j}$ is an appropriate substitution for the initial token $y_{i}$, given the context. The context is the input sequence $\bm{x}$ and the output sequence up to the token that is to be substituted $\bm{y}_{<i}$. Substituting with a very unlikely token might undermine the coherence of the overall sequence. Thus, our third objective is to favor alternative tokens that exhibit a high likelihood. The alternative token’s importance score
$I_{ij}\$$=\ p(v_{j}\mid\bm{y}_{<i},\bm{x},\bm{w})$ (11)
@@
In summary, the three scores work together to identify the appropriate initial and alternative token pair that should be substituted to alter the semantics of the initial output sequence. The attribution score identifies which initial token to substitute, while the substitution score identifies which alternative token to replace the initial token with. The importance score identifies which alternative token the language model considers a fit for substitution. Each of the scores positively contributes to the performance of SDLG, as detailed in Sec. E of the appendix.
@@
0: Language model $g(\cdot)$, input sequence $\bm{x}$, vocabulary $\mathcal{V}$, number output sequences NN
@@
[…]
@@
Figure 2: Algorithm: Illustration of the workflow of SDLG, sampling diverse yet coherent output sequences by identifying and substituting semantically important tokens in the initial output sequence using three distinct context-sensitive scores.
@@
It has to be noted that substituting initial tokens not corresponding to the beginning of a word is often impractical. Consequently, we exclusively apply substitutions to tokens at the beginning of a word, making them even more efficient. All token pairs ($y_{i}$, $v_{j}$) are ranked according to the three scores ($A_{i}$, $S_{ij}$, $I_{ij}$). In its simplest form, the ranking is based on equally weighting the three scores, which we found to work well empirically. Based on this token score ranking, a new output sequence is generated by deliberately substituting the highest-ranked token pair. Subsequent tokens are discarded as they are conditioned on the substituted token, which affects their likelihood. The remainder of the new output sequence is then generated by the language model using the usual sampling strategy (see Alg. 1 and 2). SDLG preserves the semantically less relevant part of the output sequence, eliminating the need to regenerate it and focusing on the part with a high chance of altering the semantic meaning. As a result, the generation process becomes computationally more efficient than sampling from scratch each time and potentially sampling multiple duplicate output sequences.
@@
Finally, we can discuss the exact form of the proposal distribution $q(\bm{y}\mid\bm{x},\bm{w})$ in Eq. (8), which is induced by SDLG. We define the proposal distribution as
@@
where $p(y_{t}\mid\bm{y}_{<t},\bm{x},\bm{w})$ is the likelihood of the token that is exchanged by SDLG. Intuitively, it means that we have to adjust the MC estimate in Eq. (6), because SDLG interferes in sampling and changes the token $y_{t}$ deterministically. For more details on the assumptions and a step-by-step derivation, see Sec. B in the appendix. Our experiments show that sampling according to this proposal distribution improves uncertainty estimation in NLG.